News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Monday 20 December 1999 7.30pm EST

Over the weekend, Chandra completed a long observation of Lambda And and observed the gamma-ray burst Target Of Opportunity with ACIS-S/HETG starting at 11:58pm 12/17. The observatory passed through the radiation belts with the associated ACIS CTI measurements and performed observations of Cas A with HRC-I, Cyg X-3 with ACIS-S/HETG, and NGC 404, A1795, A2841, NGC 4494, NGC 4725, and NGC 4594 with ACIS-S.

A sequence of real-time activities were conducted as planned on Monday 12/20 between 11:30am EST and 5pm [7.30pm; rjb] EST. A software patch to implement continuous dither was uplinked/tested and backed-out, and a test was started to operate the ACIS focal plane at -125 deg C for a 7 hour period and clock the CCD's. All real-time activities were completed sucessfully. SCS 107 (SI Safing) was triggered during the continuous dither test bright star hold test and executed nominally. This SCS should have been disabled during the test and a check is underway to identify why this was not seen during ASVT testing. Analysis will be completed of the continuous dither test data before final uplink in the new year. The ACIS focal plan will be commanded to return to -110 deg C at 1am EST 12/21.

During the next 48 hours Chandra will pass through the radiation belts with associated ACIS CTI measurements and continue with observations of Cen A, RX J0822-4300 and a long observation of PSR J0218+4232 with HRC-I.

The spacecraft systems continued to operate nominally.





Return to Status Reports